OpenStudy (anonymous):

a pet store sells aquariums that are rectangular prisms.the volume V of an aquarium is given by the formula V=L.W.H where L is the length,w is the width and h is the height. solve the formula for H. W=30inch. L=13inch.

OpenStudy (whpalmer4):

\[V=L*W*H\] W=30 inches L = 13 inches \[V=L*W*H\]divide both sides by L*W \[\frac{V}{L*W}=\frac{L*W*H}{L*W}\]cancel out matching L&W on right side\[\frac{V}{L*W}=H\]Substitute given values of L&W\[H=\frac{V}{30 in*13 in}=\frac{V}{390 {in}^2}\] So, given V (in in^3), we can find H (in inches) by dividing V by 390.

OpenStudy (anonymous):

thank you! so H=390?

OpenStudy (whpalmer4):

No, H = V/390, if V is in cubic inches (and H will be in inches). Think of it this way: L*W is the area of the bottom of the tank. Multiplying that by H stacks up layers to create the volume. Similarly, if we've already got the volume, dividing by the area gives us the number of those layers.

OpenStudy (anonymous):

???im lost so you need the number of the volume to get the height?

OpenStudy (anonymous):

cuz the volume is 5850 but i need the height so how does it work?